The weight of a group of women has a population mean, u, of 121 pounds, and a population standard deviation, o, of 16 pounds. The distribution is right-skewed. Suppose a random sample is taken of 100 of these women's weights.
The expected mean weight is about 121 pounds because the sample is an unbiased estimator of the population mean.
The actual sample will not be exactly the value determined above. The amount it typically differs from this value is given by the standard error. What is the standard error for a sample mean taken from this population in pounds?
Thank you!!

Answer by rothauserc(4718)

a distribution that is skewed to the right is bunched up toward the left and with a "tail" stretching toward the right.
the sample mean is the same as the population mean
standard error for a sample mean = (standard deviation * sample mean) = standard deviation / square root (N) where N is the sample size
standard error for the sample mean is standard deviation / square root (N) = 16 / square root(100) = 16/10 = 8/5 = 1.6 pounds.
It is not possible to calculate the median which is a better choice than the mean in a skewed distribution.